OpenStudy (anonymous):

how do i factor: 14x^2-47x+7=0

OpenStudy (anonymous):

You don't.

OpenStudy (anonymous):

oh.. ok well my original problem i needed to solve was -42x^2+141x=-21 and i thought i was on the right track... maybe not?

OpenStudy (anonymous):

find delta and factor using the solutions

OpenStudy (anonymous):

what is delta?

OpenStudy (anonymous):

The only thing that will make 14x^2 is (7x )*(2x ). Now, the only factors that give the 7 is 7 and 1. With this knowledge you get (7x-1)*(2x-7)

OpenStudy (anonymous):

Wrong; is it possible that taking out 3 was all you were meant to do? Because it can;t go any further.

OpenStudy (anonymous):

hmm.. maybe but the example i followed came up with two terms which then i put to the zero factor property... but maybe this is as far as this one goes.

OpenStudy (anonymous):

what about the p and q thing?

OpenStudy (anonymous):

I put it through the answer check and it said that 14x^2-47x+7 was wrong

OpenStudy (anonymous):

it said the methods to solving were either the quadratic formula or by completing the square.

OpenStudy (anonymous):

\[\sqrt{47^2-4 \times 7 \times 14} \notin \mathbb{Z^+}\] therefore it won't factorise Completing the square doesn't look too friendly on this one; just go with the formula

OpenStudy (anonymous):

\[x = \frac{47 \pm \sqrt{1817}}{28}\] ...dunoo how that will go in your 'answer check'

OpenStudy (anonymous):

haha yeah.. probably not that well.. but thanks for the help @INewton

OpenStudy (anonymous):

just an update... apparently 14x^2-47x+7=) factors into (7x+1)(2x-7)... which in turn would make the real number solutions: -1/7 and 7/2